OpenStudy (anonymous):

y = -2 x 2 (squared) + 12 x -13 =

OpenStudy (jiteshmeghwal9):

\[-2 \times (2)^2+12 \times -13\]use PEMDAS\[-2 \times 4 +12 \times -13\]\[-8+(-156)\]\[-8-156=?\]